OpenStudy (anonymous):

Find the Solution of this equation -4x-8=16

OpenStudy (anonymous):

-4x-8=16 +8 +8 -4x=24 divide by -4 what do you get?

OpenStudy (anonymous):

Move the -8 to the RHS. 24 divided by -4 is -6.

OpenStudy (anonymous):

-6